Run 12 timeline

From RHIC Spin Group
Jump to navigationJump to search

Last modified by Dmitri Smirnov on 22-05-2012


Run 11 timeline Run 12 timeline Run 13 timeline


Jan 17, 2012
RHIC setup started
Jan 30, 2012
p-Carbon polarimeters setup complete. Preliminary calibration established
Feb 08, 2012 - fill 16386
First good fill
Feb 10, 2012 @ ~1617 - fill 16400
Physics declared
Feb 18, 2012
To mitigate the noise effect the event trigger threshold is increased in Y1 and other polarimeters
Feb 22, 2012
APEX session
Mar 9, 2012
Test runs for stochastic cooling studies
In Y1D disabled detector 3 at 24 GeV. First fill 16533, 16534
Mar 14, 2012
Big access day. 42/48 targets replaced. Stochastic cooling cable terminated
Mar 15, 2012
First measurements at 255 GeV
Mar 17, 2012
Running with longitudinal beam polarisation for the experiments. First fill 16582
Apr 11, 2012
All targets replaced
Replaced grids for B1 detector 4, B2 detectors 3 and 4, and Y2 detector 4
All 45 degree grids in B2 were found pushed ~2 cm towards detectors. Fixed
As of 9:30 Yellow targets installed
Apr 18-19, 2012
H-jet injection study fill 16737


Last modified by Dmitri Smirnov on 22-05-2012